# Urban Region Representation Learning: A Flexible Approach

### v1

@@ Appendix D Details of Grid-based Inter-view Feature Learning
Appendix E Details of Grid-based Dual-feature Attentive Fusion
We employ the dual-feature attentive fusion module (DAFusion) in HAFusion (Sun et al., 2024) to generate the final cell embeddings $\mathbf{E}$ based on the cell embeddings $\mathbf{Z}$. DAFusion consists of two sub-modules: view-aware attentive fusion (ViewFusion) and region-aware attentive fusion (We call it cell-aware attentive fusion, denoted as CellFusion, since we are using this module at the cell level).
E.1. View-aware Attentive Fusion.
ViewFusion leverages the attention mechanism to learn fusion weights of the views to aggregate the view-based cell embeddings $\{\mathbf{Z}^{p},\mathbf{Z}^{l},\mathbf{Z}^{gn},\mathbf{Z}^{si}\}$. It first computes correlation scores between different views as follows:
@@ Appendix E Details of Grid-based Dual-feature Attentive Fusion
E.2. Cell-aware Attentive Fusion.
CellFusion further applies self-attention on the embeddings $\widetilde{\mathbf{Z}}$ learned by ViewFusion, to encode the higher order correlations among the learned cell embeddings. Embeddings $\widetilde{\mathbf{Z}}$ are first fed into a self-attention module to produce the hidden representations $\hat{\mathbf{Z}}$ of the grid cells (which resemble Equations 15 and 16).
@@
Appendix F Details of Multi-task Learning Objectives
Given the learned cell embeddings $\mathbf{E}$, we employ MLP to obtain four feature-oriented embeddings $\mathbf{E}^{p}$, $\mathbf{E}^{l}$, $\mathbf{E}^{n}$, and $\mathbf{E}^{si}$, which are used with the following objective functions.
F.1. POI-Oriented Objective
We utilize the graph reconstruction task to reconstruct the POI adjacent matrix $\mathbf{A}^{p}$ using the POI task embeddings $\mathbf{E}^{p}=\{e^{p}_{i}\}^{m}_{i=1}$. The objective function $\mathcal{L}^{p}$ is formulated as follows:
@@ F.1. POI-Oriented Objective
F.2. Land Use-Oriented Objective
The land use objective mirrors the POI objective, aiming to reconstruct the land use adjacent matrix $\mathbf{A}^{l}$ using the land use-oriented embeddings $\mathbf{E}^{l}$. The objective function $\mathcal{L}^{p}$ is defined as follows:
@@ F.2. Land Use-Oriented Objective
F.3. Geographic Neighbor-Oriented Objective
By the First Law of Geography (Tobler, 1970), nearby grid cells are likely to have similar functionality and embeddings. Accordingly, we utilize the triplet loss (Schroff et al., 2015) as the geographic neighbor objective function $\mathcal{L}^{gn}$ based on $\mathbf{E}^{gn}$. This loss aims to minimize the distance between an anchor and its positive sample while maximizing the distance between the anchor and a negative sample, ensuring a margin of separation. Formally, given the task embeddings $\mathbf{E}^{gn}$, the geographic neighbor objective function is formulated as:
@@ F.3. Geographic Neighbor-Oriented Objective
F.4. Satellite Image-Oriented Objective
We leverage the object counting task to predict the total number of POIs within the satellite image of each grid cell. Satellite images contain POI information, such as buildings, roads, farmlands, and shops. This task encourages the model to extract POI features from the satellite images.
@@ F.4. Satellite Image-Oriented Objective
Appendix G Cell Textual Description Example
We provide an example of the generated textual description for a hexagonal grid cell in Fig. 9. We collect POI information from OpenStreetMap and generate detailed street address using the reverse geocoding functionality of the Nominatim API (Nominatim, 2024).
@@ Appendix G Cell Textual Description Example
Appendix H Details of Experimental Datasets
@@
We use real data from five cities across the globe: New York City (NYC(New York Dataset, 2020), Chicago (CHI(Chicago Dataset, 2020), San Francisco (SF(San Francisco Dataset, 2020), Singapore City (SG(Singapore Dataset, 2019), and Lisbon (LX(Lisbon Region Boundary, 2024). We collect data on region division, POI, land use, satellite images, street view images, check-in, and population. Additionally, crime and service call data are collected for NYC, CHI, and SF (these are unavailable for the other two cities).
Table 5 summarizes the NYC, CHI, and SF datasets. Each dataset consists of region boundaries obtained from open data portals, POIs extracted from OpenStreetMap (OpenStreetMap, 2024) with category labels detailed in Appendix C, and land use data also sourced from OpenStreetMap. Satellite images are collected from Google Maps (Google Maps Platform, 2024) at a fixed resolution of 800×800 pixels, while street view images, also from Google Maps, have a resolution of 640×500 pixels and cover sampled points distributed across the full area of interest (the sampling strategy is detailed in Appendix C). Crime and service call records are retrieved from open data portals, while check-in records are obtained from a Foursquare dataset (Foursquare Dataset, 2015). Each record contains location and time information, with counts aggregated at region level. Population data for 2020 is sourced from WorldPop (WorldPop, 2024).
@@
Appendix I Details of Baseline Models
We compare against the following baseline models, categorized into two groups: those using readily accessible data and those using human mobility data.
Models based on readily accessible data:
  • CityFM (Balsebre et al., 2024) (SOTA) leverages geospatial entities (e.g., buildings and road segments) extracted from OpenStreetMap and employs contrastive learning with three objectives to generate entity embeddings: a mutual information-based text-to-text objective, a vision-language objective, and a road-based context-to-context objective. The resulting entity embeddings are then aggregated to form corresponding region embeddings.
Models based on human mobility data:
  • MGFN (Wu et al., 2022) constructs multiple mobility graphs based on the human mobility features. It then clusters these graphs into seven groups according to their spatio-temporal distances and sums the graphs within each group to form seven mobility pattern graphs. Finally, message passing is performed on the mobility pattern graphs to generate the region embeddings.
  • HREP (Zhou et al., 2023) uses human mobility, POI, and geographic neighbour features to generate region embeddings. Subsequently, it randomly generates learnable prompt embeddings and concatenates them with the region embeddings, customizing the embeddings for different downstream tasks.
  • ReCP (Li et al., 2024) uses human mobility and POI features to generate region embeddings via a multi-view learning approach. Instead of fusing multi-view information in a posteriors stage, it does the integeration by employing two objective functions: maximizing mutual information between views and minimizing conditional entropy
  • HAFusion (Sun et al., 2024) (SOTA) uses human mobility, POI, and land use features to generate region embeddings through a multi-view learning approach. It employs an attention-based fusion module to fuse multi-view information both at the region level and the view level, to capture higher-order correlations among the regions.
Table 8. Prediction Accuracy over Cities in Different Countries
Singapore Check-in Population
MAE $\downarrow$ RMSE $\downarrow$ $R^{2}\uparrow$ MAE $\downarrow$ RMSE $\downarrow$ $R^{2}\uparrow$
RegionDCL 1588 $\pm$ 38 2443 $\pm$ 45 0.102 $\pm$ 0.033 5894 $\pm$ 170 7979 $\pm$ 220 0.408 $\pm$ 0.034
UrbanCLIP 1404 $\pm$ 11 2396 $\pm$ 31 0.136 $\pm$ 0.022 6477 $\pm$   58 8651 $\pm$   82 0.305 $\pm$ 0.008
CityFM 1345 $\pm$ 40 2251 $\pm$ 59 0.239 $\pm$ 0.039 5834 $\pm$   41 7948 $\pm$   64 0.412 $\pm$ 0.010
FlexiReg 1246 $\pm$ 24 2143 $\pm$ 41 0.309 $\pm$ 0.027 4845 $\pm$ 125 6712 $\pm$ 214 0.581 $\pm$ 0.027
Improvement 7.4% 4.8% 29.3% 17.0% 15.6% 41.0%
Lisbon Check-in Population
MAE $\downarrow$ RMSE $\downarrow$ $R^{2}\uparrow$ MAE $\downarrow$ RMSE $\downarrow$ $R^{2}\uparrow$
RegionDCL 260.1 $\pm$ 22.3 474.4 $\pm$ 24.9 0.457 $\pm$ 0.056 4003 $\pm$ 301 6657 $\pm$ 313 0.586 $\pm$ 0.039
UrbanCLIP 251.2 $\pm$   3.8 411.8 $\pm$   6.1 0.591 $\pm$ 0.012 3274 $\pm$   51 4618 $\pm$   74 0.801 $\pm$ 0.004
CityFM 286.5 $\pm$   3.6 578.3 $\pm$   5.4 0.195 $\pm$ 0.004 3547 $\pm$   10 6321 $\pm$   25 0.627 $\pm$ 0.003
FlexiReg 205.6 $\pm$ 13.9 264.2 $\pm$ 17.8 0.831 $\pm$ 0.022 1844 $\pm$   56 2654 $\pm$ 171 0.934 $\pm$ 0.008
Improvement 18.2% 35.8% 40.6% 43.7% 42.5% 16.6%
Appendix J Model Hyperparameter Settings
All models were trained and tested on a machine equipped with an NVIDIA Tesla V100 GPU and 64 GB of memory.
The region embedding dimensionality $d$ is set as 144 for our model following HAFusion and HREP. The region embedding dimensionalities for MVURE, MGFN, RegionDCL, and ReCP are 96, 96, 64, and 96, respectively, as suggested by their original papers. Our experimental results in Appendix M.4 also show that these dimensionality values are optimal for the respective models (i.e., their yielded embeddings are of lower quality when $d$ is 144). The region embedding dimensionalities of UrbanCLIP and CityFM are difficult to change from their default implementation. For UrbanCLIP, $d=768$ which is determined by the image encoder, as it uses the embeddings of satellite images from a vision language model CLIP (Radford et al., 2021) as the region embedding. Similarly, for CityFM, $d=1792$ which is determined by the dimensionality of the embeddings of different geospatial entities enclosed by a region.
Appendix K Full Results on Overall Model Comparison
Table 7 presents the complete results on an overall comparison of FlexiReg against the baseline models, on the three US cities NYC, CHI, and SF for four downstream tasks, evaluated with three performance metrics MAE, RMSE, and $R^{2}$. It can be seen that FlexiReg achieves the best results consistently across all three datasets, four downstream tasks, and three evaluation metrics.
Appendix L Full Results on Cross-country Applicability
Table 7 presents the complete experimental results on the two cities SG and LX outside the USA. As it shows, FlexiReg also outperforms all competitors consistently, indicating the strong applicability of FlexiReg to cities in different countries.
Appendix M Additional Experimental Results
This section reports additional experimental results on model running time, model applicability to suburban areas, the impact of the design choice of the grid cell structure, and the impact of model parameter values.
M.1. Model Running Time.
@@ Appendix I Details of Baseline Models
Table 9 reports the running times for embedding learning and downstream task learning (we omit the times on SG and LX for conciseness as the comparative patterns resemble). The downstream task running times include both model training (prompt learning) and inference, with inference times indicated in parentheses. The downstream tasks share identical input and output sizes.
Our model FlexiReg requires additional time to learn the region embeddings because it starts with learning embeddings for the cells, and there are more cells than regions. Additionally, our model takes additional time for downstream task training, as it integrates textual features and street view visual features into the region embeddings through two alignment modules. Adding these times together, our embedding learning can still be down within less than 10 minutes, which is rather affordable for training a model with deep learning. While FlexiReg takes extra learning times, it significantly reduces prediction errors for the downstream tasks, as shown in the experiments above.
@@
The models that use readily accessible data, UrbanCLIP and CityFM, take more times for embedding learning than ours. This is because of large training datasets and complex model structures. These two models also take extra time at the training stage for the downstream tasks as they need to extract useful information from embedding by using deep networks with multiple layers. RegionDCL is slow on CHI, because the number of buildings is extremely high compared to other cities. The other baseline models, which use human mobility data, are typically faster in embedding learning, as human mobility data provides critical insights into regional interactions and the functional relationships between regions, enabling the model to converge quickly.
@@
We note that the inference times for the downstream tasks are very similar across all models, as all prediction models share a simple structure.
M.2. Applicability to Suburban Areas

### v3

@@ 3.2.2. Grid-based Inter-view Feature Learning
3.2.3. Grid-based Dual-Feature Attentive Fusion
We adopt the dual-feature attentive fusion module (DAFusion) from HAFusion (Sun et al., 2024) to generate the final cell embeddings $\mathbf{E}$ based on the cell embeddings $\mathbf{Z}$. DAFusion consists of two sub-modules: view-aware attentive fusion (ViewFusion) and region-aware attentive fusion (We call it cell-aware attentive fusion, denoted as CellFusion, since we are using this module at the cell level).
View-aware Attentive Fusion. ViewFusion leverages the attention mechanism to learn fusion weights of the views to aggregate the view-based cell embeddings $\{\mathbf{Z}^{p},\mathbf{Z}^{l},\mathbf{Z}^{gn},\mathbf{Z}^{si}\}$. It first computes correlation scores between different views as follows:
@@ 3.2.3. Grid-based Dual-Feature Attentive Fusion
Cell-aware Attentive Fusion. CellFusion further applies self-attention on the embeddings $\widetilde{\mathbf{Z}}$ learned by ViewFusion, to encode the higher order correlations among the learned cell embeddings. Embeddings $\widetilde{\mathbf{Z}}$ are first fed into a self-attention module to produce the hidden representations $\hat{\mathbf{Z}}$ of the grid cells (which resemble Equations 4 and 5).
@@
3.2.4. Module Training.
POI-Oriented Objective We utilize the graph reconstruction task to reconstruct the POI adjacent matrix $\mathbf{A}^{p}$ using the POI task embeddings $\mathbf{E}^{p}=\{e^{p}_{i}\}^{m}_{i=1}$. The objective function $\mathcal{L}^{p}$ is formulated as follows:
@@
Land Use-Oriented Objective The land use objective mirrors the POI objective, aiming to reconstruct the land use adjacent matrix $\mathbf{A}^{l}$ using the land use-oriented embeddings $\mathbf{E}^{l}$. The objective function $\mathcal{L}^{p}$ is defined as follows:
@@
Geographic Neighbor-Oriented Objective By the First Law of Geography (Tobler, 1970), nearby grid cells are likely to have similar functionality and embeddings. Accordingly, we utilize the triplet loss (Schroff et al., 2015) as the geographic neighbor objective function $\mathcal{L}^{gn}$ based on $\mathbf{E}^{gn}$. This loss aims to minimize the distance between an anchor and its positive sample while maximizing the distance between the anchor and a negative sample, ensuring a margin of separation. Formally, given the task embeddings $\mathbf{E}^{gn}$, the geographic neighbor objective function is formulated as:
@@
Satellite Image-Oriented Objective We leverage the object counting task to predict the total number of POIs within the satellite image of each grid cell. Satellite images contain POI information, such as buildings, roads, farmlands, and shops. This task encourages the model to extract POI features from the satellite images.
@@
3.3. Adaptive Region Embedding Generation
Here, $\mathbf{e}_{i}\in\mathbf{E}$ is the cell embedding of $c_{i}$.
3.4. Prompt Enhancer for Task Learning
3.4.1. Text-Region Encoding
An example of the generated textual description of a cell is shown in Fig. 4. We collect POI information from OpenStreetMap and generate detailed street address using the reverse geocoding functionality of the Nominatim API (Nominatim, 2024).
@@ 3.4.1. Text-Region Encoding
We use the same cell-to-region embedding aggregation approach as described in Section 3.3 to obtain the region embeddings from textual features, i.e., following Equations 20 and  21. The textual embeddings of regions are denoted as $\mathbf{H}^{t}\in\mathbb{R}^{n\times d_{llm}}$.
Figure 5. Text-region alignment module
3.4.2. Street View-Region Encoding
(24) $\mathbf{u}_{i,j}=\mathrm{ResNet}(\mathbf{sv}_{i,j}).$
where $|\mathbf{sv}_{i}|$ is the number of images associated with $c_{i}$.
Figure 6. Street view image embedding learning
3.4.3. Model Training.
where $y_{i}$ is the ground truth for $r_{i}$, and $n$ is the number of regions.
4. Experiments
We run experiments to verify: (Q1) the embedding quality of our FlexiReg model as compared with the state-of-the-art (SOTA) models on four downstream tasks, (Q2) the applicability of FlexiReg across diverse geographic regions, (Q3) the adaptability of our cell embeddings to different region formations, (Q4) the impact of our model components and input features, (Q5) the applicability of FlexiReg to areas of different urban environments (Q6) the impact of the grid cell design, prompt templates, and key hyper-parameters.
4.1. Experimental Settings
Dataset. We use real data from five cities across the globe: New York City (NYC(New York Dataset, 2020), Chicago (CHI(Chicago Dataset, 2020), San Francisco (SF(San Francisco Dataset, 2020), Singapore City (SG(Singapore Dataset, 2019), and Lisbon (LX(Lisbon Region Boundary, 2024). We collect data on region division, POI, land use, satellite images, street view images, check-in, and population. Additionally, crime and service call data are collected for NYC, CHI, and SF (these are unavailable for the other two cities).
@@ 4.1. Experimental Settings
Table 3 summarizes the NYC, CHI, and SF datasets. Each dataset consists of region boundaries obtained from open data portals, POIs extracted from OpenStreetMap (OpenStreetMap, 2024) with category labels detailed in Section 3.1.2, and land use data also sourced from OpenStreetMap. Satellite images are collected from Google Maps (Google Maps Platform, 2024) at a fixed resolution of 800×800 pixels, while street view images have a resolution of 640×500 pixels and cover sampled points distributed across the full area of interest (the sampling strategy is detailed in Section 3.1.2). Crime and service call records are retrieved from open data portals, while check-in records are obtained from a Foursquare dataset (Foursquare Dataset, 2015). Each record contains location and time information, with counts aggregated at region level. Population data for 2020 is sourced from WorldPop (WorldPop, 2024).
@@
Competitors. We compare with models from two categories. The first category uses a subset of the publicly accessible data: RegionDCL (Li et al., 2023), UrbanCLIP (Yan et al., 2024), and CityFM (Balsebre et al., 2024) (SOTA). The second category uses human mobility data, which has restricted availability: MVURE (Zhang et al., 2020), MGFN (Wu et al., 2022), HREP (Zhou et al., 2023), ReCP (Li et al., 2024), and HAFusion (Sun et al., 2024) (SOTA). This latter category does not apply to SG and LX due to lack of data.
Models based on readily accessible data:
  • CityFM (Balsebre et al., 2024) (SOTA) leverages geospatial entities (e.g., buildings and road segments) extracted from OpenStreetMap and employs contrastive learning with three objectives to generate entity embeddings: a mutual information-based text-to-text objective, a vision-language objective, and a road-based context-to-context objective. The resulting entity embeddings are then aggregated to generate corresponding region embeddings.
Models based on human mobility data:
  • MGFN (Wu et al., 2022) constructs multiple mobility graphs based on the human mobility features. It then clusters these graphs to form mobility pattern graphs based on their spatio-temporal distances. Finally, message passing is performed on the mobility pattern graphs to generate the region embeddings.
  • HREP (Zhou et al., 2023) uses human mobility, POI, and geographic neighbour features to generate region embeddings. Subsequently, it randomly generates learnable prompt embeddings and concatenates them with the region embeddings, tailoring the embeddings for different downstream tasks.
  • ReCP (Li et al., 2024) uses human mobility and POI features to generate region embeddings via a multi-view learning approach. Instead of fusing multi-view information in a posterior stage, it does the integration by employing two objective functions: maximizing mutual information between views and minimizing conditional entropy.
  • HAFusion (Sun et al., 2024) (SOTA) uses human mobility, POI, and land use features to generate region embeddings. It employs an attention-based fusion module to fuse multi-view information both at the region level and the view level, to capture higher-order correlations among the regions.
Model hyperparameter settings. All models were trained and tested on a machine equipped with an NVIDIA Tesla V100 GPU and 64 GB of memory.
The region embedding dimensionality $d$ is set as 144 for our model following HAFusion and HREP. The region embedding dimensionalities for MVURE, MGFN, RegionDCL, and ReCP are 96, 96, 64, and 96, respectively, as suggested by their original papers. Our experimental results in Section 4.10.1 also show that these dimensionality values are optimal for the respective models (i.e., their yielded embeddings are of lower quality when $d$ is 144). The region embedding dimensionalities of UrbanCLIP and CityFM are difficult to change from their default implementation. For UrbanCLIP, $d=768$ which is determined by the image encoder, as it uses the embeddings of satellite images from a vision language model CLIP (Radford et al., 2021) as the region embedding. Similarly, for CityFM, $d=1792$ which is determined by the dimensionality of the embeddings of different geospatial entities enclosed by a region.
We evaluate the representation learning models through the downstream models in mean absolute error (MAE), root mean square error (RMSE), and coefficient of determination ($\boldsymbol{R^{2}}$).
4.2. Overall Results (Q1)
The overall model accuracy results are reported in Table 5. We make the following observations.
@@ 4.2. Overall Results (Q1)
Model running time. Table 6 reports the running times for embedding learning and downstream task learning (we omit the times on SG and LX as the comparative patterns resemble). The downstream task running times include both model training (prompt learning) and inference, with inference times in parentheses. The downstream tasks share identical input and output sizes.
Our model FlexiReg requires additional time to learn the region embeddings because it starts with learning embeddings for the cells, and there are more cells than regions. Additionally, our model takes additional time for downstream task training, as it integrates textual features and street view visual features into the region embeddings through two alignment modules. Adding these times together, our embedding learning can still be done in less than 10 minutes, which is rather affordable for training a model with deep learning. While FlexiReg takes extra learning times, it significantly reduces prediction errors, as shown in the experiments above.
@@
The models that use readily accessible data, UrbanCLIP and CityFM, take more times for embedding learning than ours. This is because of large training datasets and complex model structures. These two models also take extra time at the training stage for the downstream tasks as they need to extract useful information from embedding by using deep networks with multiple layers. RegionDCL is slow on CHI, because the number of buildings is extremely high compared to other cities. The other baseline models, which use human mobility data, are typically faster in embedding learning 00, as human mobility data provides critical insights into regional interactions and the functional relationships between regions, enabling the model to converge quickly.
@@
The inference times for the downstream tasks are similar across all models, as all prediction models share a simple structure.
Table 7. Impact of ##\##Cells to Embedding Learning Model Running Time (in seconds)
##\##Cell 128 256 512 1024 2048
Training time 73 125 228 445 803
Inference time 0.055 0.105 0.204 0.412 0.746
We further evaluat model running times as the number of cells varies from 128 to 2048 (cells are our basic embedding unit). As shown Table 7, the running times increase linearly with the number of cells. With 2048 cells, training takes about 10 minutes, and inference takes less than 1 second (on an NVIDIA Tesla V100 GPU). These results validate our model scalability.
4.3. Cross-country Applicability (Q2)
Table 8. Prediction Accuracy over Cities in Different Countries
Singapore Lisbon
Check-in Population Check-in Population
$R^{2}\uparrow$ $R^{2}\uparrow$ $R^{2}\uparrow$ $R^{2}\uparrow$
CityFM 0.239 $\pm$ 0.039 0.412 $\pm$ 0.010 0.195 $\pm$ 0.004 0.627 $\pm$ 0.003
Improvement 29.3% 41.0% 40.6% 16.6%
Table 8 reports the results. We only report $R^{2}$ for conciseness, as the performance in MAE and RMSE resembles (Same below). FlexiReg outperforms all competitors consistently, with an improvement of at least 16.6% in $R^{2}$ over the best baseline models. These results confirm FlexiReg’s applicability across countries. Among the baseline models, CityFM performs better on Singapore, while UrbanCLIP is more suitable for Lisbon, underscoring the different urban characteristics of the two cities.
4.4. Adaptability to Region Formations (Q3)
(a) Crime
(b) Check-in
(c) Service Call
(d) Population
Figure 7. Adaptability to regions of varying sizes (NYC).
4.5. Ablation Study (Q4)
We study the effectiveness of FlexiReg model components with the following variants: (1) FlexiReg-w/o-PE removes the prompt enhancer from the downstream task learning stage. (2) FlexiReg-w/o-TAlign replaces the text-region alignment module with a direct concatenation of text and region embeddings. (3) FlexiReg-w/o-SVAlign replaces the street view-region alignment module with the summation of all street view image embeddings in a region, followed by concatenating these with the region embeddings. (4) FlexiReg-w/o-EC removes the environment context-based contrastive learning and directly use ResNet to generate visual embeddings for each street view image. (5) FlexiReg-w/o-CNN replaces the CNN branch with a GNN branch to process satellite images. (6) FlexiReg-w/o-Grid directly learns region embeddings without using the grid cells. (7) FlexiReg-w/o-FS uses all features (including textual descriptions and street view images) during cell embedding learning. (8) FlexiReg-w/o-WS replaces the weighted summation of cell embeddings with a direct summation. (9) FlexiReg-w/o-LT uses the average of all token embeddings as the text embeddings instead of the last token embeddings.
Figure 8. Ablation study results (NYC).
4.6. Impact of Input Features (Q4)
Figure 9. Impact of input features (NYC).
Regarding the issue of street view image availability, FlexiReg-w/o-SV simulates scenarios where such imagery is unavailable for all regions. The results demonstrate that our model continues to achieve competitive performance by leveraging the remaining features and model components. Additionally, we perform a more fine-grained experiment to simulate partial availability, where 25% of regions lack street view imagery, and we randomly sample street view images from their neighboring regions. As shown in Table 9, the model’s performance remains largely unaffected, indicating its robustness to incomplete street view data.
Table 9. The availability of street view images ($R^{2}\uparrow$ on NYC)
Crime Check-in Service Call Population
100% 0.789 0.876 0.601 0.715
75% 0.785 0.870 0.595 0.711
4.7. Applicability to Suburban Areas (Q5)